OpenStudy (anonymous):

Algebra Help What is the equation of the line that has a slope of 3 and goes through point (4, 7)?

OpenStudy (whpalmer4):

Okay, there's a formula you should know called point-slope form. For a line with slope \(m\) going through a point \((x_0,y_0)\): \[y-y_0 =m(x-x_0)\]

OpenStudy (anonymous):

I know the formula but need help solving

OpenStudy (whpalmer4):

to be a bit pedantic, there are an infinite number of equations that have that slope and go through that point :-) For example, if \[y = 3x\]is the equation (it isn't), then \[y+1 = 3x+1\]\[y+0.5=3x+0.5\]\[y-3x=0\]\[-3x=-y\]etc. are all equations that go through the same points with the same slope.

OpenStudy (whpalmer4):

Okay, you have \(x_0 = 4, y_0 = 7, m = 3\) plug them into the formula, what do you get?

OpenStudy (whpalmer4):

I'll do the left hand side: \[y-y_0 = m(x-x_0)\]\[y-7 = m(x-x_0)\]Now you do the rest

OpenStudy (anonymous):

I'm Lost

OpenStudy (whpalmer4):

Copy "\(y-7 = m(x-x_0)\)" on to your paper. As you write, if you encounter an \(m\), replace it with "3". If you encounter an \(x_0\), replace it with "4". What do you get?